# Circumgalactic Enrichment by a Multi-Galaxy Merger at Redshift 6.7

### v1

@@ Abstract
Recent JWST observations have uncovered high-redshift galaxies characterized by multiple star-forming clumps, many of which appear to be undergoing mergers [1, 2, 3, 4]. Such mergers are powerful drivers of galaxy evolution, triggering intense star formation [5, 6, 7] that accelerates the formation of massive quiescent galaxies [8]. Mergers can also expel and redistribute the metal-enriched gas from galaxies into their circum-galactic medium (CGM), regulating the chemical component of their environments [9, 10, 11] and creating channels of ionized gas through which ionizing radiation escapes [12, 13]. Here, we report a merger of at least five galaxies, dubbed JWST’s Quintet (JQ), at redshift 6.7, discovered in the JWST GOODS-South field. This system resides in a small area ( pkpc2), containing over 17 galaxy-size clumps with a total stellar mass of . The JQ system has a total star formation rate of 240 – 270 yr-1, placing it dex above the median star formation rate–mass main sequence at this epoch. The high mass and star formation rate of the JQ galaxies are sufficient to produce massive, inactive galaxies identified at later times, by redshift 4 [14, 15, 16], suggesting this may be a progenitor event for such systems. We detect a large [O III]+H emitting gaseous halo surrounding and connecting four galaxies in the JQ, consistent with the interaction between them. This indicates that the CGM of JQ is already enriched by the heavy elements generated from the galaxies through merger-induced tidal stripping or galactic outflows. Our findings offer a plausible evolutionary pathway for the formation of massive quiescent galaxies observed at lower redshifts and provide direct evidence for the metal enrichment of their environments through galaxy mergers, just 800 Myr after the Big Bang.
@@
We identified a group of five emission line galaxies (ELGs) at redshift 6.7, which we call JWST’s Quintet (JQ), from a study of ELGs based on imaging and spectroscopic data from the JWST Advanced Deep Extragalactic Survey [17] (JADES). JADES, one of the deepest JWST/NIRCam imaging surveys to date, provides continuous wavelength coverage from 0.8 to 5.0 m across 14 broad- and medium-band filters (see Methods). The JADES multi-wavelength imaging data enables precise photometric redshift estimation with an uncertainty of [17], allowing reliable distances to galaxies and distinguishing galaxies in close physical proximity. We identified ELGs with strong [O III]+H nebular emission as well-detected () sources with photometric redshifts in the range 6.7 to 7.6 with a excess of flux emission in the F410M medium-band image compared to the F444W broad-band image of JADES. At these redshifts, the [O III] and H emission lines fall in the F410M bandpass, and the flux excess criteria identities objects with EW([O III]+H) 300 Å. The spatial distribution of ELGs revealed an overdensity of the five JQ galaxies within a region, see Figure 1, corresponding to proper kpc2 (pkpc2) Of these, two galaxies (ELG1 and ELG5) have spectroscopically confirmed redshifts of 6.707 and 6.712 based on the NIRSpec spectroscopy [18]. These two galaxies reside at the head and the tail of the JQ with a projected distance of 18.6 pkpc and a velocity difference of 195 km s-1. The remaining three ELGs (ELG2, ELG3, ELG4) reside between them and have similar spectral energy distributions (SEDs) as ELG1 and ELG5 with consistent photometric redshifts (see Figure 2). This is evidence all five ELGs are in the same physical structure.
@@
The JQ appears to be a high-redshift analog of a well-known galaxy group, Stephan’s Quintet, at a distance of 85 Mpc. Though, unlike Stephan’s Quintet, where only four of the five galaxies are physically associated and interacting in mergers [19, 20], we believe all galaxies in JQ are physically associated. Furthermore, there is evidence that the entire structure of ELG1, ELG2, ELG3, and ELG4 is connected by a diffuse emission of [O III]+H only seen in the F356W image. At the redshift of JQ, the [O III] line falls within the F410M band, facilitating the detection of this system, but the F356W band encompasses both [O III] and H with much higher transmission than F410M. We therefore use F356W for analysis in this paper. The implication is that these galaxies have undergone an interaction that produced the diffuse [O III]+H emission. In ELG2, a notably extended tail (labeled C5 in Figure 1) points toward ELG1 in the projected plane, with several faint star-forming clumps (C4, C12, C13; clump identification is described below) distributed along it. These clumps likely trace the star formation induced by the galaxy mergers. ELG3 also shows a disturbed morphology with two clumps (C8-1, C8-2) connected through diffuse emission to ELG4. While we do not detect diffuse [O III]+H emission nor star-forming clumps between ELG5 and ELG3, the small velocity offset of 195 km s-1 between ELG5 and ELG1 and the small projected distance of 6.2 pkpc between ELG5 and ELG3 strongly suggests that ELG5 is also interacting with the other ELGs.
@@
The JQ galaxies show a high frequency of spatially resolved clumps. We adopt an automatic clump identification algorithm combined with visual inspection to detect the clumps in this structure. We measure the ELG sizes and photometry by fitting their surface brightness with Sérsic profile models, and we measure the galaxy properties by fitting models to the SEDs. We identified 17 clumps associated with the JQ and 7 foreground galaxies. Extended Data Table 1 lists the properties of the JQ clumps. To account for the diffuse emission surrounding the clumps, we also perform a spatially resolved SED fitting to measure the integrated properties of the entire system. We derive a total stellar mass of and star formation rate (SFR) of yr-1. We remove the foreground galaxies from the images based on their best-fit Sérsic profiles and present a clean map of JQ in Figure 1c,d. See Methods for details.
@@
The sizes and SFRs of the JQ clumps are significantly larger than those of the star clusters and star cluster-size clumps within individual galaxies resolved through the lensing technique [21, 22, 23]. Their sizes, stellar masses, and SFRs are more consistent with the high-redshift spatially unresolved field galaxies observed with JWST NIRCam [24]. Although recent JWST observations have reported numerous galaxy systems containing such galaxy-size clumps [4, 2, 3, 1], JQ stands out as the first system identified with over 10 galaxy-size clumps. This high density of galaxy-size clumps within a single galaxy system is extremely rare and challenges predictions from cosmological simulations. For example, work focusing on clump formation during mergers between high-redshift galaxies does not identify any systems with more than four clumps at [25]. Comparing JQ to predictions from the 6.5 – 7.0 galaxy systems in a 2-deg2 simulation [26], we identify only 8 systems with galaxy-size clumps out from a parent sample of 29,696 systems. Furthermore, from this simulation, we would expect only 0.26 JQ-like systems in the JADES survey at – 7.6. This indicates the number density of galaxy-size clumps in the JQ is higher than galaxy systems in cosmological simulation, and is a rare configuration. The small area of JADES F410M image ( arcmin2) implies an observed comoving number density of cMpc3 for JQ, almost quadruple the expected number density from the cosmological simulation ( cMpc3). see Methods for details.
@@
The clumps associated with the JQ galaxies show evidence of elevated SFRs compared to other galaxies at these epochs. Figure 3 compares the SFRs of the clumps to the galaxies at similar redshift and the SFR–Mass Main Sequence (SFMS) using results from the literature [27, 29, 30, 28]. In all cases, the clumps in JQ with stellar masses have SFRs more than 1 dex above the measured SFMS relation, indicating they are in a bursting phase [27]. The best-fit model star-formation histories of these clumps show that they could form all of their stellar mass in the previous 10 Myr, but this is a lower limit as the recently formed stars could outshine previous generations of stars [32, 33]. Nevertheless, the short formation timescales suggest that the star formation in these clumps was recently triggered by major mergers in JQ [25]. The clumps with higher stellar mass, (C3, C8, C9) have SFRs consistent with the median SFR–Mass relation at this epoch. The total SFR of galaxies in the JQ system is higher than the SFMS by dex. This indicates that the elevated SFR and the formation of clumps is correlated with the major merger in the JQ system.
@@
The elevated SFR triggered by major mergers in JQ provides a plausible evolutionary pathway for forming massive quiescent galaxies, recently discovered at 4.5 – 5. The existence of these objects has challenged our understanding of rapid galaxy growth and quenching over short timescales [14, 16, 31]. The high stellar mass and SFR of the JQ system are consistent with those needed to produce massive (), quiescent galaxies observed at redshift 4 – 5 [14, 16] (see Figure 4). The expected merging timescale for the JQ system is – 350 Myr (see Methods). If the JQ galaxies sustain the current SFR during this period, they will build a single massive galaxy with a stellar mass of SFR – . Simulations show that such a major merger event is expected to rapidly consume cold gas and trigger strong feedback from black hole activities, ultimately quenching star formation in the JQ galaxies [34, 35, 36]. Furthermore, the observed number density of JQ-like systems is consistent with that of massive quiescent galaxies at redshift , estimated to be Mpc-3 [16, 14]. These factors suggest that JQ is a likely candidate for a progenitor of the massive quiescent galaxies identified at redshifts 4 – 5. To test this interpretation requires a detailed study of the kinematics of the individual JQ clumps, which is possible with future JWST observations.
@@
The NIRCam F356W image of JQ is dominated by the redshifted [O III]+H emission line. The image shows extended, diffuse emission, compared to the NIRCam F335M and F277W images, which is evidence of a [O III]+H halo surrounding the JQ system. We estimate the amount of [O III]+H emission by subtracting the continuum contribution from the F356W image (see Methods) and present the [O III]+H halo as the orange contours in Figure 1 (the original [O III]+H map is shown in Extended Data Figure 3). The [O III]+H halo surrounds the entirety of the ELG1, ELG2, ELG3, and ELG4 galaxies, extending more than 20 kpc. The total [O III]+H luminosity is erg s-1. Notably, we observe a “bridge” in the [O III]+H emission connecting galaxies ELG1 and ELG2. This bridge extends to 5 kpc and includes the clumps C12 and C13. The presence of this large-scale [O III]+H halo requires oxygen-enriched gas in the circum-galactic medium (CGM), ejected and/or stripped efficiently from the galaxies in JQ, particularly from ELG1 and ELG2, as evidenced by the bright [O III]+H bridge connecting them.
@@
It is therefore more plausible that the [O III]+H halo of JQ results from oxygen-enriched gas stripped out of the galaxies through interactions and tidal forces [11, 41]. The stripped gases are shock-heated in galaxy collisions [7, 42], and dispersed to large radii leading to large hot gaseous halos [43]. This scenario is believed to be responsible for the diffuse emission halo in Stephan’s Quintet, where the shocked-heated gas has a temperature of K and appears as a ridge of X-ray and radio synchrotron emission [44, 45], residing between the two colliding galaxies. Such emission is below the sensitivity of X-ray and radio telescopes given JQ’s distance, but it is possible to confirm the shock-heated origin of the gas using excitation diagnostics of emission lines, as those have been performed in Stephan’s Quintet [46, 47, 48]. Similar to the distribution of shock-heated gas in Stephan’s Quintet, the majority of diffuse [O III]+H emission in JQ also arises from the “bridge” between two merging galaxies, ELG1 and ELG2. Therefore, the evidence supports the interpretation that the [O III]+H halo in JQ results from shock-heated gas, stripped from the galaxies through interactions from a merger.
The discovery of the JQ system at in the JADES data suggests that major mergers of multiple galaxies might be more frequent in the early universe than expected from theory. These mergers can efficiently eject metal-enriched gas into the CGM, thereby contributing to the development of the galactic ecosystem. The elevated star formation associated with mergers offers a unique solution to the tension between the observation and cosmological simulation on the observed overabundance of the massive galaxies and massive quiescent galaxies recently unveiled by JWST.
Methods
@@
HST and JWST Imaging data
We utilize the published JWST images and spectroscopy from the JWST Advanced Deep Extragalactic Survey (JADES) DR3 [50, 51, 18] to measure the flux densities and to detect clumps in the JQ galaxies. JADES is one of the deepest NIRCam surveys, providing 14 broad- or medium-bands (broad- and medium-band imaging from F090W, F115W, F150W, F182M, F200W, F210M, F277W, F335M, F356W, F410M, and F444W filters from JADES [2], combined with medium-band imaging in F430M, F460M, and F480M [52]) covering wavelengths from 0.8 to 5.0 m. The JADES photometric catalog provides matched-aperture flux densities for the galaxies, and photometric redshifts measured from the SEDs [17, 51, 2].

### v2

# Extended Enriched Gas in a Multi-Galaxy Merger at Redshift 6.7

@@ Abstract
Recent JWST observations have uncovered high-redshift galaxies characterized by multiple star-forming clumps, many of which appear to be undergoing mergers [1, 2, 3, 4]. Such mergers, especially those of two galaxies with equivalent masses, play a critical role in driving galaxy evolution [5, 6, 7, 8, 9] and regulating the chemical composition of their environments [10, 11]. Here, we report a major merger of at least five galaxies, dubbed JWST’s Quintet (JQ), at redshift 6.7, discovered in the JWST GOODS-South field. This system resides in a small area ( pkpc2), containing over 17 galaxy-size clumps with a total stellar mass of . The JQ system has a total star formation rate of 240 – 270 yr-1, placing it dex above the median star formation rate–mass main sequence at this epoch. The high mass and star formation rate of the JQ galaxies are consistent with the star formation history of those unexpected massive quiescent galaxies observed at redshift 4 – 5 [12, 13, 14], offering a plausible evolutionary pathway for the formation of such galaxies. We also detect a large [O III]+H emitting gaseous halo surrounding and connecting four galaxies in the JQ, suggesting the existence of heavy elements in the surrounding medium–inner part of its circumgalactic medium (CGM). This provides direct evidence for the metal enrichment of galaxies’ environments through merger-induced tidal stripping, just 800 Myr after the Big Bang.
@@
We identified a group of five emission line galaxies (ELGs) at redshift 6.7, which we call JWST’s Quintet (JQ), from a study of ELGs based on imaging and spectroscopic data from the JWST Advanced Deep Extragalactic Survey [15] (JADES). JADES, one of the deepest JWST/NIRCam imaging surveys to date, provides continuous wavelength coverage from 0.8 to 5.0 m across 14 broad- and medium-band filters (see Methods). The JADES multi-wavelength imaging data enables precise photometric redshift estimation with an uncertainty of [15], allowing reliable distances to galaxies and distinguishing galaxies in close physical proximity. We identified ELGs with strong [O III]+H nebular emission as well-detected () sources with photometric redshifts in the range 6.7 to 7.6 with a excess of flux emission in the F410M medium-band image compared to the F444W broad-band image of JADES. At these redshifts, the [O III] and H emission lines fall in the F410M bandpass, and the flux excess criteria identify objects with EW([O III]+H) 300 Å. The spatial distribution of ELGs revealed an overdensity of the five galaxies within a region, see Figure 1, corresponding to proper kpc2 (pkpc2) Of these, two galaxies (ELG1 and ELG5) have spectroscopically confirmed redshifts of 6.707 and 6.712 based on the NIRSpec spectroscopy [16]. These two galaxies reside at the head and the tail of the JQ with a projected distance of 18.6 pkpc and a velocity difference of 195 km s-1. The remaining three ELGs (ELG2, ELG3, ELG4) reside between them and have similar spectral energy distributions (SEDs) as ELG1 and ELG5 with consistent photometric redshifts (see Figure 2). This is evidence that all five ELGs are in the same physical structure.
@@
The JQ appears to be a high-redshift situation similar to that in the well-known galaxy group, Stephan’s Quintet, at a distance of 85 Mpc. Unlike Stephan’s Quintet, where only four of the five galaxies are physically associated and interacting in mergers [17, 18], we believe all galaxies in JQ are physically associated. Furthermore, there is evidence that the entire structure of ELG1, ELG2, ELG3, and ELG4 is connected by a diffuse emission of [O III]+H only seen in the F356W image. At the redshift of JQ, the [O III] line falls within the F410M band, facilitating the detection of this system, but the F356W band encompasses both [O III] and H with much higher transmission than F410M. We therefore use F356W for analysis in this paper. The implication is that these galaxies have undergone an interaction that produced the diffuse [O III]+H emission. In ELG2, a notably extended tail (labeled C5 in Figure 1) points toward ELG1 in the projected plane, with several faint star-forming clumps (C4, C12, C13; clump identification is described below) distributed along it. These clumps likely trace the star formation induced by the galaxy mergers. ELG3 also shows a disturbed morphology with two clumps (C8-1, C8-2) connected through diffuse emission to ELG4. The velocity offset (195 km s-1) between ELG5 and ELG1, and the projected distance (6.2 pkpc) between ELG5 and ELG3, are both smaller than the virial velocities and radii of these ELGs (see Methods), suggesting that ELG5 is gravitationally bound to the JQ system. Furthermore, ref [19] reported an unusually high Ly escape fraction () in ELG5, suggesting it resides within a large ionized bubble – too large to be produced by ELG5 alone. Generating such an ionized bubble requires substantial contributions from the entire JQ system. These facts support that ELG5 is associated with the JQ. However, we caution that the degeneracy between line-of-sight distance and velocity offset introduces uncertainty; it remains possible that the true three-dimensional separation between ELG5 and the other ELGs exceeds the virial radius, in which case ELG5 would not yet be bound.
@@
The JQ galaxies show a high frequency of spatially resolved clumps. We adopt an automatic clump identification algorithm combined with visual inspection to detect the clumps in this structure. We measure the clump sizes and photometry by fitting their surface brightness with Sérsic profile models, and we measure the galaxy properties by fitting models to the SEDs. We identified 17 clumps associated with the JQ and 7 foreground galaxies. Extended Data Table 1 lists the properties of the JQ clumps. To account for the diffuse emission surrounding the clumps, we also perform a spatially resolved SED fitting to measure the integrated properties of the entire system. We derive a total stellar mass of and star formation rate (SFR) of yr-1. We remove the foreground galaxies from the images based on their best-fit Sérsic profiles and present a clean map of JQ in Figure 1c,d. See Methods for details.
@@
The sizes and SFRs of the JQ clumps are significantly larger than those of the star clusters and star cluster-size clumps within individual galaxies resolved through the lensing technique [20, 21, 22]. Their sizes, stellar masses, and SFRs are more consistent with the high-redshift spatially unresolved field galaxies observed with JWST NIRCam [23]. We, therefore, define them as galaxy-size clumps. Although recent JWST observations have reported numerous galaxy systems containing such galaxy-size clumps [3, 2, 24, 1], JQ stands out as the first system identified with over 10 galaxy-size clumps. This high density of galaxy-size clumps within a single galaxy system is extremely rare and challenges predictions from cosmological simulations. For example, work focusing on clump formation during mergers between high-redshift galaxies does not identify any systems with more than four clumps at [25]. Comparing JQ to predictions from the 6.5 – 7.0 galaxy systems in a 2-deg2 simulation [26], we identify only 8 systems with galaxy-size clumps out of a parent sample of 29,696 systems , and 4 of which are classified as major mergers. From this simulation, we would expect only 0.13 JQ-like systems in the JADES survey at – 7.6. This indicates the number density of galaxy-size clumps in the JQ is higher than galaxy systems in cosmological simulation, and is a rare configuration. The small area of JADES F410M image ( arcmin2) implies an observed comoving number density of cMpc3 for JQ, almost one dex larger than the expected number density from the cosmological simulation ( cMpc3). See Methods for details.
@@
The five galaxies and their associated clumps show evidence of elevated SFRs compared to typical galaxies at these epochs. Figure 3 compares the SFRs of the clumps to the galaxies at similar redshift and the SFR–Mass Main Sequence (SFMS) using results from the literature [27, 29, 30, 28]. All five galaxies and clumps, except ELG3 and C16, have SFRs more than 1 dex above the observed SFMS relation [27], and the total SFR of JQ is also higher than the SFMS by dex, indicating they are in a bursting phase. Best-fit star-formation histories show that these galaxies and clumps could form all of their stellar mass in the past 10 Myr, but this is a lower limit, as the recently formed stars could outshine previous generations of stars [32, 33]. Nevertheless, the short formation timescales imply that the star formation in these galaxies and clumps was recently triggered by major mergers in JQ [25]. In contrast to those less massive components, the most massive galaxy, ELG3 with a stellar mass of , contributing of the total stellar mass, shows no evidence of an enhanced SFR. Its SFR accounts for of total SFR and lies on the median SFMS relation at this epoch. This suggests that this merger has impacted the star formation of the lower-mass galaxies and clumps more than ELG3. Consequently, the future mass growth of JQ will be driven by the combination of its star formation along the main sequence and its merging companions, consistent with the ex-situ star formation scenario suggested for massive major mergers [34].
@@
The substantial ex-situ star formation and elevated SFR triggered by major mergers in JQ provides a plausible evolutionary pathway for forming massive quiescent galaxies, recently discovered at 4.5 – 5. The existence of these objects has challenged our understanding of rapid galaxy growth and quenching over short timescales [12, 14, 31, 35]. The high stellar mass and SFR of the JQ system are consistent with those needed to produce massive (), quiescent galaxies observed at redshift 4 – 5 [12, 14] (see Figure 4). The expected merging timescale for the JQ system is – 350 Myr (see Methods). If the JQ galaxies sustain the current SFR during this period, they will build a single massive galaxy with a stellar mass of SFR – . Simulations show that such a major merger event is expected to rapidly consume cold gas and trigger strong feedback from black hole activities, ultimately quenching star formation in the JQ galaxies [36, 37, 38]. Furthermore, the observed number density of JQ-like systems is consistent with that of massive quiescent galaxies at redshift , estimated to be Mpc-3 [14, 12]. These factors suggest that JQ is a likely candidate for a progenitor of the massive quiescent galaxies identified at redshifts 4 – 5. To test this interpretation requires a detailed study of the kinematics of the individual JQ clumps. This can be achived with future observations from JWST NIRSpec integral field unit or NIRCam wide field slitless spectroscopy (WFSS). In particular, a NIRCam WFSS observation using the F356W filter is expected in fall 2025 [39], which will provide the kinematics and [O III]+H fluxes of bright clumps in JQ.
@@
The NIRCam F356W image of JQ is dominated by the redshifted [O III]+H emission line. The image shows extended, diffuse emission, compared to the NIRCam F335M and F277W images, which is evidence of a [O III]+H halo surrounding the JQ system. We estimate the amount of [O III]+H emission by subtracting the continuum contribution from the F356W image (see Methods) and present the [O III]+H halo as the green contours in Figure 1 (the original [O III]+H map is shown in Extended Data Figure 4). The [O III]+H halo surrounds the entirety of the ELG1, ELG2, ELG3, and ELG4 galaxies, extending more than 20 kpc. The total [O III]+H luminosity is erg s-1. Notably, we observe a “bridge” in the [O III]+H emission connecting galaxies ELG1 and ELG2. This bridge extends to 5 kpc and includes the clumps C12 and C13. The presence of this large-scale [O III]+H halo requires oxygen-enriched gases in the surrounding medium of galaxies, the inner part of the CGM. These oxygen-enriched gases were ejected and/or stripped efficiently from the galaxies in JQ, particularly from ELG1 and ELG2, as evidenced by the bright [O III]+H bridge connecting them.
@@
It is therefore more plausible that the [O III]+H halo of JQ results from oxygen-enriched gas stripped out of the galaxies through interactions and tidal forces [44, 11]. The stripped gases are shock-heated in galaxy collisions [6, 45], and dispersed to large radii, leading to large hot gaseous halos [46]. This scenario is believed to be responsible for the diffuse emission halo in Stephan’s Quintet, where the shocked-heated gas has a temperature of K and appears as a ridge of X-ray and radio synchrotron emission [47, 48], residing between the two colliding galaxies. Such emission is below the sensitivity of X-ray and radio telescopes given JQ’s distance, but it is possible to confirm the shock-heated origin of the gas using excitation diagnostics of emission lines, as those have been performed in Stephan’s Quintet [49, 50, 51]. Similar to the distribution of shock-heated gas in Stephan’s Quintet, the majority of diffuse [O III]+H emission in JQ also arises from the “bridge” between two merging galaxies, ELG1 and ELG2. Therefore, the evidence supports the interpretation that the [O III]+H halo in JQ results from shock-heated gas, stripped from the galaxies through interactions from a merger.
The discovery of the JQ system at in the JADES data suggests that major mergers of multiple galaxies might be more frequent in the early universe than expected from theory. These mergers can efficiently eject metal-enriched gas into the CGM, thereby contributing to the development of the galactic ecosystem. The elevated star formation associated with mergers offers a unique solution to the tension between the observation and cosmological simulation on the observed overabundance of massive galaxies and massive quiescent galaxies recently unveiled by JWST.
1 Methods
@@
We acknowledge that in this work, we use the term “clump” to describe a compact, spatially unresolved structure in the JWST image. We define a galaxy as an isolated system composed of one or more clumps with similar stellar populations; we do not impose a specific minimum number of clumps for this definition. An emission line galaxy is defined as a galaxy showing strong equivalent widths in nebular lines, such as [O III]+H.
1.1 HST and JWST Imaging data
We utilize the published JWST images and spectroscopy from the JWST Advanced Deep Extragalactic Survey (JADES) DR3 [53, 54, 16] to measure the flux densities and to detect clumps in the JQ galaxies. JADES is one of the deepest NIRCam surveys, providing 14 broad- or medium-bands (broad- and medium-band imaging from F090W, F115W, F150W, F182M, etc. filters from JADES [2], combined with medium-band imaging in F430M, F460M, and F480M from JEMS [55]) covering wavelengths from 0.8 to 5.0 m. The JADES photometric catalog provides matched-aperture flux densities for the galaxies, and photometric redshifts measured from the SEDs [15, 54, 2].